OpenStudy (anonymous):

Find the rate of change of the distance between the origin and a moving point on the graph of y=x^2+1 if dx/dt=7 centimeters per second

OpenStudy (amistre64):

use the distance formula, derive it, and plug in the values

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

@amistre64 I would love to see you solve it.

OpenStudy (amistre64):

d^2 = x^2 + y^2 dd' = xx' + yy', x' = 7

OpenStudy (anonymous):

and x and y are equal to zero?

OpenStudy (amistre64):

so,\[d' = \frac{7x+yy'}{\sqrt{x^2+y^2}}\] y' can be determined from the given setup for y

OpenStudy (anonymous):

What values do we plug in for x and y?

OpenStudy (amistre64):

whatever values satisfy the definition for y of course.

OpenStudy (amistre64):

the rate of change is a function of time, they dont give you a specific point to asses so you have to let it be assessed for any point that satisifies the setup of course

OpenStudy (anonymous):

but if we plug in (0,0) the answer would be all zero.

OpenStudy (amistre64):

y=x^2+1 0,0 is not a point of concern

OpenStudy (amistre64):

now, if we are creating a vector field, then yeah we would run into that situation, but in this case the domain is not RxR, but a subset of it that does not include 0,0

OpenStudy (anonymous):

okay can you solve it please then? what values do you plug in for x and y? You derive them from the equation?

OpenStudy (amistre64):

x is independant, y is already given x' is stated, and y' can be solved d' is defined by the derivative of the distance formula ... which part is not working out?

OpenStudy (anonymous):

So after you derived the formula, what values do you plug in for x and y? This is my question.

OpenStudy (amistre64):

The subset of RxR (the ordered pairs) that satisfy y = x^2 + 1 as is stated in the problem

OpenStudy (amistre64):

\[D=\{(x,y)\in RxR:y=x^2+1\}\]

OpenStudy (amistre64):

i spose \cross would be more appealing, or the lazier R^2

OpenStudy (anonymous):

okay so for y we plug in the formula and for x you plug in the x?

OpenStudy (amistre64):

yep

OpenStudy (amistre64):

and y' = 2xx' = 14x

OpenStudy (anonymous):

@amistre64 Thank you. I had a lot of trouble with these types of questions too. I really appreciate your help.

OpenStudy (amistre64):

its all about reading the information, and not being blinded by the whatifs :)

OpenStudy (amistre64):

dd' = xx' + yy', x' = 7, y' = 2xx' = 14x \[d' = \frac{7x+14x}{\sqrt{x^2+(x^2+1)^2}}\]

OpenStudy (amistre64):

opps forgot the y up top

OpenStudy (amistre64):

\[d' = \frac{7x+14x(x^2+1)}{\sqrt{x^2+(x^2+1)^2}}\] thats better

OpenStudy (anonymous):

@lamp12345 btw its a really nice question I really like it.

OpenStudy (amistre64):

its a thinker, doable but coy

OpenStudy (anonymous):

@amistre64 I was understanding it too but thats not an answer choice. Here they are

OpenStudy (amistre64):

simplify is needed of course. i figure that by the time your doing this stuff that you already are familiar and sufficient with your algebra

OpenStudy (amistre64):

also, if you could put all the pics in one file it would do better than trying to hunt and peck thru 5 seperate files

OpenStudy (anonymous):

yeah sorry I typed that out before you gave the other answer

OpenStudy (amistre64):

when i do implicits i like to keep my x' in there since its a place holder for dx/dt. IF this had been a dx/dx setup then x'=1, but dx/dt = 7 is their definition

OpenStudy (amistre64):

for the double chk http://www.wolframalpha.com/input/?i=+simplify+%5Cfrac%7B7x%2B14x%28x%5E2%2B1%29%7D%7B%5Csqrt%7Bx%5E2%2B%28x%5E2%2B1%29%5E2%7D%7D 14x^3 + 21x ------------ sqrt(.....) seems to fit